I need help with this problem: b^4a - 81a^5
Answer by Alan3354(69443)

I need help with this problem: b^4a - 81a^5
-------------------
Help doing what? Maybe you want to factor it?
b^4a - 81a^5
= a(b^4 - 81a^4)
= a(b^2 - 9a^2)*(b^2 + 9a^2)
= a(b - 3a)*(b + 3a)*(b^2 + 9a^2)
If you need help doing something else, you'll have to be more specific.